O is the centre of a circle HKL. |HK|=16cm, |HL|=10cm and the perpendicular from O to HK is 4cm. What is the length of the perpendicular from O to HL?
Answer by addingup(3677)

There is no HKL circle, a circle does not have sides or corners that you can name. Is HKL a triangle? Or are HK and KL chords on a circle?
------------------------
IF chords on a circle, first we find the radius:
16/2 = 8 we have two triangles with long leg 8 and short leg 4. The hypotenuse of these triangles is our radius:
sqrt(8^2+4^2) = 8.94
--------------------------------
Now we find distance to HL:
10/2 = 5 this is the long leg of our two triangles. We know the hypotenuse, it's our radius 8.94. We have to find the other leg, the distance from O to HL. Pythagoras comes to the rescue again:
sqrt(8.94^2-5^2) = 7.41 this is the distance from the center to HL